A 14 m long truck is driven alongside a 500 m long train with the same velocity of

25 m/s. The front of the truck is 7 m behind the rear of the train. At a certain instant, the

truck stars accelerating at 0,08 m/s2, while the train starts decelerating at 0,4 m/s2. How

long will it take before the rear of the truck passes the front of the train and what distance

would the truck have covered in this time?


1
Expert's answer
2021-08-16T02:33:01-0400

1) After t sec the truck passes the front of the train

"S = ut+\frac{1}{2}at^2"

Initial velocity "u=0"

"493 = 0+\frac {1}{2}\times 0.08 t^2 = 111"  sec

2) Let the distance be d meters

When they met

"60+d= 25\times111-\frac{1}{2} \times0.4\times111^2"

"d = 310.8-60=250.8" m